### v2

# GABAR: Graph Attention-Based Action Ranking for Relational Policy Learning

Abstract
We propose a novel approach to learn relational policies for classical planning based on learning to rank actions. We introduce a new graph representation that explicitly captures action information and propose a Graph Neural Network architecture augmented with Gated Recurrent Units (GRUs) to learn action rankings. Our model is trained on small problem instances and generalizes to significantly larger instances where traditional planning becomes computationally expensive. Experimental results across standard planning benchmarks demonstrate that our action-ranking approach achieves generalization to significantly larger problems than those used in training.
Introduction
Classical planning tackles the problem of finding action sequences to achieve goals in deterministic environments. While traditional planners can solve small problems optimally using search and heuristics, they often struggle with scalability. This has motivated research into learning general relational policies from small solved problems, which can then be applied to significantly larger instances (Ståhlberg, Bonet, and Geffner 2022a). The key insight behind learning-based planning is that optimal solutions to small problems often reveal patterns that generalize to bigger problems. For example, in Blocks World, solutions to $4$ block problems can teach a policy to stack blocks bottom-up that generalizes to problems with $20$ or more blocks. The advantage of relational learning is its ability to capture compositional structure. This compositionality also enables strong generalization (Zambaldi et al. 2018), (Fern, Yoon, and Givan 2003) (Džeroski, De Raedt, and Driessens 2001).
Learning approaches in automated planning have traditionally focused on learning heuristic functions to guide search algorithms. These methods typically learn a value function that estimates the distance to the goal and integrate it within classical search algorithms like A* or greedy best-first search. The learned heuristics help focus the search but still require explicit search during plan execution.
Recently, several neural approaches have been proposed for learning domain-specific heuristics. (Toyer et al. 2020) learn neural heuristics that generalize across problem instances but still rely on search for planning. Similarly, (Shen, Trevizan, and Thiébaux 2020) use hypergraph networks to learn heuristics that can transfer across problems. Other approaches that learn value functions for planning include (Chen, Thiébaux, and Trevizan 2023), (Chen, Thiébaux, and Trevizan 2024). While these approaches show promise in learning useful search guidance, they suffer from the inherent computational overhead of the search process during planning.
An alternative line of work focuses on learning policies that can directly select actions without search. Value-based methods like (Ståhlberg, Bonet, and Geffner 2022b), (Ståhlberg, Bonet, and Geffner 2022a) learn a value function that induces a greedy policy by selecting actions leading to states with minimum estimated cost-to-go. While effective, these approaches face two key limitations: First, value functions are not only quite complex and challenging to learn, but also are unnecessary for action selection. Second, in domains where optimal planning is NP-hard (like the Blocks World), optimal value functions are not easily generalizable to larger problem sizes (Gupta and Nau 1992).
Learning to Rank approaches have shown promise in planning domains but have primarily focused on ranking states rather than actions. (Garrett, Kaelbling, and Lozano-Pérez 2016) pioneered this direction by using RankSVM to learn state rankings using hand-crafted features. More recently, (Chrestien et al. 2024) and (Hao et al. 2024) demonstrated that learning to rank states can be more effective than learning precise heuristic values, as the relative ordering of states is sufficient for guiding search. However, these approaches still fundamentally rely on search during execution, inheriting signficant computational overhead.
Our work is inspired by the effectiveness of graph neural networks (GNNs) to represent and learn general relational policies such as (Ståhlberg, Bonet, and Geffner 2022b, 2023, 2024; Chen, Thiébaux, and Trevizan 2024). We introduce a novel architecture GABAR (Graph Attention-Based Action Ranking), which directly learns to rank actions rather than estimating value functions. GABAR introduces (1) an action-centric graph representation that explicitly captures how objects participate in actions and (2) a GNN architecture with Gated Recurrent Units (GRUs) that learns how object representations update based on their role in actions. One of our key insights is that ranking actions that are applicable in the same state often turns out to be easier and more generalizable than ranking states by their distances to the goals. Through experiments on standard benchmarks, we show GABAR achieves generalization to significantly larger problems than those used for training.
Problem Setup
Classical planning deals with finding a sequence of actions that transform an initial state into a goal state. A classical planning problem is represented as a pair $P=\langle D,I\rangle$, where $D$ represents a first-order domain and $I$ contains instance-specific information. The domain $D$ consists of a set of predicate symbols $\mathcal{P}$ with associated arities and a set of action schemas $\mathcal{A}$. Each action schema $a\in\mathcal{A}$ is defined by a set of parameters $\Delta(a)$ representing variables that can be instantiated, preconditions $pre(a)$, add effects $add(a)$, and delete effects $del(a)$. The instance information $I$ is a tuple $\langle O,s_{0},G\rangle$ where $O$ is a finite set of objects, $s_{0}$ is the initial state represented as a set of ground atoms $p(o_{1},...,o_{k})$ where $p\in\mathcal{P}$ and $o_{i}\in O$, and $G$ is the goal condition also represented as a set of ground atoms.
A state $s$ is a set of ground atoms that are true in that state. An action schema can be grounded by substituting its parameters with objects from $O$. A ground action $a$ is applicable in state $s$ if $pre(a)\subseteq s$, and results in successor state $s^{\prime}=(s\setminus del(a))\cup add(a)$. A solution or plan is a sequence of applicable ground actions that transform the initial state $s_{0}$ into a state satisfying the goal condition $G$. A relational policy maps a problem state to an action. The current paper addresses the following problem. Given a domain $D$ and a set of training instances of different sizes and their solutions, learn a relational policy that leads to efficient solutions for larger test instances from the same domain.
Graph Attention-Based Action Ranking
Figure 1: GABAR’s architecture for action extraction. (a) Graph representation: The input PDDL problem is converted into a graph with four types of nodes (predicate, object, action schema, and global) connected by predicate-object and action-object edges that encode state and grounded action information. (b) GNN encoder: Processes the graph through $L$ rounds of message passing where edge, node, and global representations are sequentially updated based on their interactions. (c) Action decoder: Uses the final global embedding to construct a grounded action through a GRU-based decoder sequentially - first selecting an action schema, then iteratively choosing objects for each parameter position until a complete grounded action is formed.
Learning general policies for classical planning domains requires effectively processing variable-sized states and selecting appropriate actions that generalize across problem instances. We introduce GABAR (Graph Attention-based Decoder for Action Ranking), a novel architecture that directly learns to rank and select actions rather than learning value functions. GABAR consists of four key components: (1) a graph-based state representation that explicitly captures grounded action information, (2) a neural encoder that processes this rich representation, and (3) a GRU-based decoder that sequentially constructs complete actions. This section details each component and explains how they work together to enable effective action selection.
System Overview
Given a planning instance, I in PDDL format (Fox and Long 2003), along with the set of ground actions, GABAR operates by first converting this to a graph structure that makes explicit the relationships between objects, predicates, and potential actions. This graph is then processed through our neural architecture to rank actions(as described in Fig 1). Then, the highest-ranked applicable action is executed to reach the next state I’. This process is repeated until a goal state is reached. To ensure the execution terminates, the system maintains a history of visited states and avoids actions that would lead to previously visited states. The execution continues until either reaching a goal state or terminating if no unvisited successor states are available or if the maximum execution length (1000 in our experiments) is exceeded.
Graph Representation
We introduce a novel graph representation for classical planning tasks that captures the structural relationships between objects, predicates, and actions and the semantic information needed for learning action ranking effectively. Our representation $G=(V,E,X,R)$ consists of a set of nodes $V$, edges $E$, node features $X$, and edge features $R$.
The node set $V=O\cup P\cup A\cup\{g\}$ where $O$, $P$,and $A$ represent sets of domain objects, grounded predicates, and action schemas respectively and $g$ is a global node that aggregates graph-level information. The edge set $E=E_{\text{pred}}\cup E_{\text{act}}$, where $E_{\text{pred}}$ is the set of edges between predicates and their argument objects and similarly $E_{\text{act}}$ is the set of edges between action schemas and their argument objects.
Node Features
The node feature function $X:V\rightarrow\mathbb{R}^{d}$ maps each node to a feature vector that encodes type and semantic information. The feature vector is constructed by concatenating several one-hot encoded segments:
For any node $v\in V$: $X(v)=[X_{\text{type}}(v)\parallel X_{\text{act}}(v)\parallel X_{\text{pred}}(v)\parallel X_{\text{obj}}(v)]$, where:
  • $X_{\text{type}}\in\{0,1\}^{3}$: One-hot encoding of node type (object, predicate, or action)
  • $X_{\text{act}}\in\{0,1\}^{|A|}$: One-hot encoding of action type (if $v$ is an action node)
  • $X_{\text{pred}}\in\{0,1\}^{2|P|}$: Encoding for predicates, where first $|P|$ bits indicate predicate type and next $|P|$ bits indicate goal predicates ((if $v$ is a predicate node))
  • $X_{\text{obj}}\in\{0,1\}^{|T|}$: One-hot encoding of object type (if $v$ is an object node), where $T$ is the set of object types
Edge Features
The edge feature function $R:E\rightarrow\mathbb{R}^{k}$ maps each edge to a feature vector encoding edge type and role information:
For any edge $e\in E$: $R(e)=[R_{\text{type}}(e)\parallel R_{\text{pred}}(e)\parallel R_{\text{act}}(e)]$, where:
  • $R_{\text{type}}\in\{0,1\}^{2}$: One-hot encoding of edge type (predicate-object or action-object)
  • $R_{\text{pred}}\in\{0,1\}^{m}$: For predicate-object edges, one-hot encoding of argument position ($m$ is max predicate arity)
  • $R_{\text{act}}\in\{0,1\}^{(m+|P|)}$: For action-object edges, concatenation of:
    • One-hot encoding of parameter position in action schema ($m$ bits)
    • Binary vector indicating which predicates are satisfied by the object in the grounded action ($|P|$ bits)
Global Features
The global node $g$ is initialized with a zero vector in $\mathbb{R}^{h}$ where $h$ is the chosen hidden representation dimension. This node can be used to aggregate and propagate graph-level information during message passing.
This graph representation captures both the structural and semantic information necessary for learning planning heuristics while maintaining a bounded feature dimension independent of problem size. The node features encode type and semantic information, while the edge features capture the relationships between objects, predicates, and actions in the planning domain.
Neural Architecture
Our neural architecture processes this graph representation through multiple components designed to handle the challenges of processing variable-sized inputs, capturing long-range dependencies between objects and actions, and making sequential decisions to construct complete actions. We detail each component below:
Graph Neural Network Encoder
Graph Neural Networks (GNNs) (Scarselli et al. 2008) are particularly well-suited for encoding states in planning problems as they can naturally process relational structures while being invariant to permutations and handling varying input sizes. This allows them to learn patterns that generalize across different problem instances within the same domain, regardless of the number of objects involved. The key insight is that planning states are inherently relational - objects interact through predicates and actions - and GNNs can capture these relationships through message passing between nodes and edges.
Our GNN encoder processes the input graph through $L$ layers of message passing between nodes, edges, and a global node. At each layer $l$, the updates proceed as follows:
First, edge embeddings are updated based on their incident nodes and the global context:
$\mathbf{e}^{l+1}_{ij}=\phi_{e}([\mathbf{e}^{l}_{ij};\mathbf{v}^{l}_{i};\mathbf{v}^{l}_{j};\mathbf{g}^{l}])$ (1)
where $\mathbf{e}^{l}_{ij}$ is the embedding of edge $(i,j)$ at layer $l$, $\mathbf{v}^{l}_{i}$ and $\mathbf{v}^{l}_{j}$ are the embeddings of its incident nodes, $\mathbf{g}^{l}$ is the global node embedding, [;][;][ ; ] denotes concatenation, and $\phi_{e}$ is a learnable neural network.
Next, node embeddings are updated using information from their connected edges and the global context:
$\mathbf{v}^{l+1}_{i}=\phi_{v}([\mathbf{v}^{l}_{i};\text{AGG}(\{\mathbf{e}^{l+1}_{ij}|j\in\mathcal{N}(i)\});\mathbf{g}^{l}])$ (2)
Where AGG is a permutation-invariant aggregation function (we use attention weighted sum), $\mathcal{N}(i)$ denotes the neighbors of node $i$, and $\phi_{h}$ is a learnable network.
Finally, the global node is updated by aggregating information from all nodes and edges:
$\mathbf{g}^{l+1}=\phi_{g}([\mathbf{g}^{l};\text{AGG}(\{\mathbf{v}^{l+1}_{i}|i\in\mathcal{V}\});\text{AGG}(\{\mathbf{e}^{l+1}_{ij}|(i,j)\in\mathcal{E}\})])$ (3)
where $\phi_{g}$ is a learnable network that combines the previous global state with aggregated node and edge information.
This architecture makes several key design choices motivated by the planning domain. First, we explicitly model and update edge representations because edges in our graph capture crucial information about action applicability. This edge information helps guide the model toward selecting valid and effective actions. Second, we include a global node that can rapidly propagate information across the graph. This is particularly important as planning problems scale up - without the global node, information would need to flow through many message-passing steps to reach distant parts of the graph. The global node acts as a shortcut, allowing the model to maintain a comprehensive view of the planning state even as the number of objects and relations grows. After $L$ rounds of message passing have been performed, the final global node embedding $g^{l+1}$ captures the relevant planning context needed for action selection.
GRU-based Action Decoder
One of the key novelties of our architecture is a GRU-based action decoder that supports actions with an arbitrary number of parameters. The decoder transforms the final global graph representation $g_{l+1}$ into a sequence of parameter decisions that fully specify an action. This sequential process is managed by a Gated Recurrent Unit (GRU) that maintains a hidden state $h_{i}$ representing the partial action construction:
$v_{1}$ $=\text{GRU}(g_{v+1},0)$ // Initialize with global state
[…]
The GRU architecture allows the decoder to maintain relevant context while making a sequence of interdependent decisions. This is crucial for ensuring that selected objects form valid and effective action instantiations.
During training, action selection proceeds greedily in two phases: First, the decoder selects an action schema using the initial hidden state:
$\text{score}(a)=\text{MLP}([v_{1};a])\quad\text{for each action schema }a$
$a^{*}=\arg\max_{a\in A}\text{softmax}(\text{score}(a))$
Then, it iteratively selects the action parameters:
$\text{score}(o)=\text{MLP}([v_{i};o])\quad\text{for each candidate object }o$
$o^{*}=\arg\max_{o\in O}\text{softmax}(\text{score}(o))$
$v_{i+1}=\text{GRU}(v_{i},\text{embed}(o^{*}))$
Since greedy parameter selection is often too myopic, we employ beam search to explore multiple choices in parallel. For a beam width $k$, at each step, we maintain the $k$ highest-scoring partial sequences. The final output is a ranked list of $k$ action groundings $(a,o_{1},\ldots,o_{n})$ along with their accumulated scores. This aligns with our goal of learning to rank actions rather than just selecting individual best choices - while training optimizes for selecting the optimal action, the learned model provides a ranking of top $k$ grounded actions during execution. The planner can then use this ranking to make more informed decisions, such as incorporating additional criteria such as cycle avoidance.
Data Generation and Training
For each planning domain, we generate training data by solving a set of small problem instances using an optimal planner. Each training example consists of a planning state $s$, goal specification $G$, and the first action $a^{*}$ from the optimal plan from $s$ to $G$. For states with multiple optimal actions, we randomly select one to avoid biasing the model.
The state-goal pairs are converted into our graph representation $G=(V,E,X,R)$ as described in the graph representation section. For each action $a^{*}$ in the training data, we create supervision signals in the form of:
  • $y_{a}$: A one-hot vector over the action schema space indicating the correct action type
  • $y_{o}=\{y_{o1},...,y_{ok}\}$: A sequence of $k$ one-hot vectors over the object space, where $k$ is the maximum number of parameters any action can take, indicating the correct objects for each parameter position
For action schema selection: the model needs to learn to assign the highest score to the correct action schema among all possible schemas. For object selection: for each parameter position, the model needs to learn to assign the highest score to the correct object among all candidate objects. This is done using the following loss function.
Loss Function Given a training instance $(G,y_{a},y_{o})$, GABAR computes action scores $s_{a}$ for all possible action schemas and object scores $s_{o_{i}}$ for each parameter position $i$. The total loss is computed as $L=L_{action}+L_{objects}$, where, $L_{objects}$ is the sum of cross-entropy losses between $\text{softmax}(s_{o_{i}})$ and $y_{o_{i}}$ for each parameter position $i$.
Training Procedure We train the model using the Adam optimizer with a learning rate of $0.0005$, $9$ rounds of GNN, and batch size of $16$. Training proceeds for a maximum of 500 epochs, and we select the model checkpoint that achieves the lowest loss on the validation set for evaluation.
Experiments
Table 1: Performance of both GABAR and GABAR-G methods across domains(number of problems shown in parenthesis(#)). Coverage indicates the percentage of test instances solved within 1000 steps. Plan lengths are reported only for solved instances.
GABAR GABAR-G
Domain Cov.(%) $\uparrow$ Mean PL $\downarrow$ Median PL $\downarrow$ PQR $\uparrow$ Cov.(%) $\uparrow$ Mean PL $\downarrow$ Median PL $\downarrow$ PQR $\uparrow$
Blocks World (100) 100 53.1 46 1.42 84 52.6 54.7 1.1
Gripper (100) 100 147.1 133 0.99 89 146.7 137 0.95
Miconic (50) 100 181.7 181 0.97 75 192.4 189 0.87
Logistics (50) 76 155.2 132 0.65 12 238.5 220 0.33
Visitall (50) 88 207.9 190 1.07 63 205.7 183 0.86
Grid (50) 96 72.1 62 0.81 68 70.2 68 0.73
We evaluate GABAR’s performance across a diverse set of classical planning domains. Our experiments aim to assess both the quality of learned policies and their ability to generalize to significantly larger problems than those in training.
Domains. We selected six standard planning domains that present different types of structural complexity and scaling dimensions.
Blocks World involves manipulating blocks to achieve specific tower configurations. The domain’s complexity scales with the number of blocks (6-9 blocks for training/validation, 10-40 blocks for testing).
Gripper requires a robot with two grippers to transport balls between rooms. The domain scales primarily with the number of balls to be moved (5-17 balls for training/validation, up to 100 balls for testing).
Miconic involves controlling an elevator to transport passengers between floors. The domain complexity increases along two dimensions: the number of passengers (1-10 for training/validation, 20-100 for testing) and the number of floors (2-20 for training/validation, 11-30 for testing).
Logistics involves transporting packages between locations using trucks (for intra-city transport) and airplanes (for inter-city transport). The domain scales with both the number of cities (4-8 for training/validation, 15-30 for testing) and packages (3-10 for training/validation, 9-24 for testing).
Visitall requires an agent to visit all cells in a grid. The domain scales with grid size (9-49 cells for training/validation, up to 400 cells for testing - testing problems 8 times larger than the training dataset).
Grid involves navigating through a grid where certain doors are locked and require specific keys to open. The number of locks (3) and keys (5) remain the same across training and testing while varying the size of the grid (7$\times$9 for training/validation to 11$\times$14 for testing - a 150% increase in cells).
Evaluation Metrics. We evaluate GABAR using three primary metrics:
  • Coverage (Cov): The percentage of test instances successfully solved within a 1000-step limit.
  • Plan Length (PL): Measured through both mean and median plan lengths across solved instances.
  • Plan Quality Ratio(PQR): Ratio of plan length produced by Fast Downward (FD) planner to the plan length produced by the learned policy. FD is run with fd-lama-first setting. We chose this satisficing configuration over optimal planners since optimal planners fail to solve most test problems within a reasonable time. While this means we cannot guarantee the optimality of the reference plans, it provides a practical baseline for assessing solution quality across our test suite.
High coverage on larger instances demonstrates the model’s ability to learn robust action selection strategies, while plan length and quality metrics reveal whether these strategies remain efficient as problems scale up.
Ablation: To understand the importance of different architectural components, we conduct an ablation study focusing on the design choice of using global nodes in the graph (GABAR-G). This variant of GABAR is without the global node, which helps assess its role in information propagation and helps quantify how the global node affects coverage, plan quality, and length.
Results and Discussions
Table 1 shows the results of our learning system across $6$ domains. We can see that the proposed method (OURS) generalizes very well on three domains - Blocksworld, Gripper, and Miconic. We can also see that our method generalizes well in the other three domains. This result, while not as good as others, shows an important capability of the proposed system - the ability to handle complex domains.
Both Grid and Logistics are more complex domains (Ståhlberg, Bonet, and Geffner 2022b), as expressing properties of the environment requires multiple relations (Ex, the feature expressing that a package is in a city while possibly within a vehicle needs multiple relations to express). We can see that the proposed method can learn these complex features that are required for solving the planning problem as well as generalize to larger problems in these settings (success rate of 76% and 96% in them). While the success rate of Visitall is not as impressive as Blocksworld, the generalization it achieves is impressive - it is solving problems up to eight times larger than what it trained on with an 88% success rate.
The ablation study comparing GABAR with and without the global node (GABAR vs. GABAR-G) reveals that removing the global node leads to degradation in performance.
In domains that primarily require local reasoning, like Blocks World, the impact is moderate but still notable - coverage drops from 100% to 84%, and the plan quality ratio decreases from 1.42 to 1.1. This suggests that while local messages passing between objects and actions can capture basic patterns, the global node helps maintain a broader context. This degradation is even more dramatic in Logistics, where coverage plummets from 76% to just 12%. The stark difference can be attributed to the need for long-range planning, as logistics requires orchestrating multiple vehicles across different cities.
Plan quality also suffers significantly without the global node. In Logistics, the plan quality ratio drops from 0.65 to 0.33, indicating that solutions become nearly twice as inefficient. This degradation in quality suggests that the global node plays a crucial role in helping the model learn strategic action selection rather than just locally reasonable choices.
These findings validate our architectural choice of including a global node in GABAR. The global node proves essential for facilitating long-range information sharing that helps coordinate actions across distant entities.
Conclusion and Future Work
We presented GABAR, a novel graph-based architecture for learning generalized policies in classical planning through action ranking. Our key contributions include (1) an action-centric graph representation that explicitly captures action-object relationships, (2) a GNN architecture augmented with global nodes and GRUs for effective information propagation, and (3) a sequential decoder that learns to construct complete grounded actions.
Our experimental results demonstrate strong generalization capabilities across multiple planning domains. GABAR achieves great coverage on Blocksworld, Gripper, Visitall, and Miconic domains when scaling to problems that are many times larger than those used in training. The architecture also shows promising results in more complex domains like Grid and Logistics. The ablation study removing global nodes highlights their critical role in enabling effective information propagation.
The primary scaling challenge lies in the graph representation’s growth rate, particularly for domains with high-arity actions or predicates. Future work could explore more compact representations while maintaining expressiveness and investigate ways for pruning irrelevant actions in the graph.
Acknowledgments
The authors acknowledge the support of Army Research Office under grant W911NF2210251

### v3

@@
-GABAR: Graph Attention-Based Action Ranking for Relational Policy Learning
+Graph Neural Network Based Action Ranking for Planning
 Abstract
 We propose a novel approach to learn relational policies for classical planning based on learning to rank actions.
-We introduce a new graph representation that explicitly captures action information and propose a Graph Neural Network architecture augmented with Gated Recurrent Units (GRUs) to learn action rankings.
-Our model is trained on small problem instances and generalizes to significantly larger instances where traditional planning becomes computationally expensive.
-Experimental results across standard planning benchmarks demonstrate that our action-ranking approach achieves generalization to significantly larger problems than those used in training.
-Introduction
+We introduce a new graph representation that explicitly captures action information and propose a Graph Neural Network (GNN) architecture augmented with Gated Recurrent Units (GRUs) to learn action rankings.
+Unlike value-function based approaches that must learn a globally consistent function, our action ranking method only needs to learn locally consistent ranking, which is more sample-efficient.
+Our model is trained on data generated from small problem instances that are easily solved by planners and is applied to significantly larger instances where planning is computationally prohibitive.
+Experimental results across standard planning benchmarks demonstrate that our action-ranking approach not only achieves better generalization to larger problems than those used in training but also outperforms multiple baseline (value function and action ranking) methods in terms of success rate and plan quality.
+1 Introduction
 Classical planning tackles the problem of finding action sequences to achieve goals in deterministic environments.
 While traditional planners can solve small problems optimally using search and heuristics, they often struggle with scalability.
-This has motivated research into learning general relational policies from small solved problems, which can then be applied to significantly larger instances (Ståhlberg, Bonet, and Geffner 2022a).
+This has motivated research into learning general relational policies from small solved problems, which can then be applied to significantly larger instances (expressive_v1, ).
 The key insight behind learning-based planning is that optimal solutions to small problems often reveal patterns that generalize to bigger problems.
 For example, in Blocks World, solutions to $4$ block problems can teach a policy to stack blocks bottom-up that generalizes to problems with $20$ or more blocks.
-The advantage of relational learning is its ability to capture compositional structure.
-This compositionality also enables strong generalization (Zambaldi et al.
-2018), (Fern, Yoon, and Givan 2003) (Džeroski, De Raedt, and Driessens 2001).
+The advantage of relational learning is its ability to capture compositional structure, which in turn enable strong generalization (relational_drl, ), (approx_Policy_iter, ) (relational_rl, ) relational_mdp_learning .
 Learning approaches in automated planning have traditionally focused on learning heuristic functions to guide search algorithms.
@@
 The learned heuristics help focus the search but still require explicit search during plan execution.
-Recently, several neural approaches have been proposed for learning domain-specific heuristics.
-(Toyer et al.
-2020) learn neural heuristics that generalize across problem instances but still rely on search for planning.
-Similarly, (Shen, Trevizan, and Thiébaux 2020) use hypergraph networks to learn heuristics that can transfer across problems.
-Other approaches that learn value functions for planning include (Chen, Thiébaux, and Trevizan 2023), (Chen, Thiébaux, and Trevizan 2024).
-While these approaches show promise in learning useful search guidance, they suffer from the inherent computational overhead of the search process during planning.
-An alternative line of work focuses on learning policies that can directly select actions without search.
-Value-based methods like (Ståhlberg, Bonet, and Geffner 2022b), (Ståhlberg, Bonet, and Geffner 2022a) learn a value function that induces a greedy policy by selecting actions leading to states with minimum estimated cost-to-go.
-While effective, these approaches face two key limitations: First, value functions are not only quite complex and challenging to learn, but also are unnecessary for action selection.
-Second, in domains where optimal planning is NP-hard (like the Blocks World), optimal value functions are not easily generalizable to larger problem sizes (Gupta and Nau 1992).
+Recently, several neural approaches have been proposed in the planning domain.
+Some methods learn value functions to guide search processes (strips_hgn, ; goose_v2, ), while others learn value functions that induce greedy policies by selecting actions leading to states with minimum estimated cost-to-go (unsupervised_gnn, ; expressive_v1, ).
+While these approaches show promise, they all face a fundamental challenge: value functions are not only complex and challenging to learn, but since optimal planning is NP-hard in most domains (complexity-planning, ; blocks_np_hard, ), there is no reason to think that optimal value functions generalize to larger problem sizes.
+In our work, instead of trying to learn optimal value functions we focus on learning a policy that ranks good actions higher in any given state.
 Learning to rank approaches have shown promise in planning domains but have primarily focused on ranking states rather than actions.
-(Garrett, Kaelbling, and Lozano-Pérez 2016) pioneered this direction by using RankSVM to learn state rankings using hand-crafted features.
-More recently, (Chrestien et al.
-2024) and (Hao et al.
-2024) demonstrated that learning to rank states can be more effective than learning precise heuristic values, as the relative ordering of states is sufficient for guiding search.
-However, these approaches still fundamentally rely on search during execution, inheriting signficant computational overhead.
-Our work is inspired by the effectiveness of graph neural networks (GNNs) to represent and learn general relational policies such as (Ståhlberg, Bonet, and Geffner 2022b, 2023, 2024; Chen, Thiébaux, and Trevizan 2024).
-We introduce a novel architecture GABAR (Graph Attention-Based Action Ranking), which directly learns to rank actions rather than estimating value functions.
-GABAR introduces (1) an action-centric graph representation that explicitly captures how objects participate in actions and (2) a GNN architecture with Gated Recurrent Units (GRUs) that learns how object representations update based on their role in actions.
-One of our key insights is that ranking actions that are applicable in the same state often turns out to be easier and more generalizable than ranking states by their distances to the goals.
-Through experiments on standard benchmarks, we show GABAR achieves generalization to significantly larger problems than those used for training.
-Problem Setup
+(ranking_og, ) pioneered this direction by using RankSVM to learn state rankings using hand-crafted features.
+More recently, (chrestien_optimize, ), and gbfs_v2 demonstrated that learning to rank states can be more effective than learning precise heuristic values, as the relative ordering of states is sufficient for guiding search.
+However, these approaches still fundamentally rely on search during execution, inheriting significant computational overhead.
+The other drawback is that, while ranking states is easier than learning the exact value function, they are still trying to learn a globally consistent state ranking function which is challenging.
+We overcome this issue in our work by learning only local ranking over actions in any given state instead of learning a ranking among all possible states.
+Unlike other action ranking approaches such as (karia2021learning, ), (rddl_gnn, ), (sr_drl, ), and (policy_gradients, ), we excitability represent action information in our input state representation.
+Our work is inspired by the effectiveness of graph neural networks (GNNs) (gnn_base, ) to represent and learn general relational policies such as (unsupervised_gnn, ; policy_gradients, ; policy_beyond_c2, ; goose_v2, ; relational_drl, ) where GNNs have been shown to be able to learn well over state representations for planning problems.
+GNNs are especially suited for these types of problems as they can handle inputs of varied sizes and can learn from graphs of small size and generalize to larger graphs hamilton2020graph .
+This is the property we are looking to exploit - train a system on small sized planning problems where it is easy to run planners and collect data and use this learned model to solve larger problems that planners are too slow to solve while simplifying the learning problem by only learning local action rankings.
+To achieve this, we introduce a novel architecture GABAR (GrAph neural network Based Action Ranking), which directly learns to rank actions rather than estimating value functions.
+GABAR consists of three key components: (1) an action-centric graph representation of state that explicitly captures how objects participate in actions (2) a GNN encoder that processes this rich representation, and (3) a GRU-based decoder that sequentially constructs parameterized actions.
+Our key insight is that ranking actions that are applicable in the same state often turns out to be easier and more generalizable than ranking states by their distances to the goals.
+Through experiments on standard benchmarks, we show that (a) GABAR achieves generalization to significantly larger problems than those used for training, and (b) it does markedly better on the larger problems when compared to value-function-based methods or other methods that do not include action information in the graph.
+2 Problem Setup
 Classical planning deals with finding a sequence of actions that transform an initial state into a goal state.
@@
 Given a domain $D$ and a set of training instances of different sizes and their solutions, learn a relational policy that leads to efficient solutions for larger test instances from the same domain.
-Graph Attention-Based Action Ranking
+3 GNN Based Action Ranking
+Learning general policies for classical planning domains requires learning to select appropriate actions effectively to variable-sized states and generalize across problem instances.
+This section details each component of GABAR, the graph representation, the GNN encoder and the GRU decoder in detail and explains how they work together to enable effective action selection.
+3.1 System Overview
+Given a planning instance, I in PDDL format pddl , along with the set of ground actions, GABAR operates by first converting this to a graph structure that makes explicit the relationships between objects, predicates, and potential actions.
+This graph is then processed through our neural architecture to rank actions(as described in Fig 1).
+Then, the highest-ranked applicable action is executed to reach the next state I’, and the process repeats.
+To ensure that the execution terminates, the system maintains a history of visited states and avoids actions that would lead to previously visited states.
+The execution continues until either a goal state is reached or a state is reached with no unvisited successor states, or the maximum execution length (1000 in our experiments) is exceeded.
 Figure 1: GABAR’s architecture for action extraction.
 (a) Graph representation: The input PDDL problem is converted into a graph with four types of nodes (predicate, object, action schema, and global) connected by predicate-object and action-object edges that encode state and grounded action information.
-(b) GNN encoder: Processes the graph through $L$ rounds of message passing where edge, node, and global representations are sequentially updated based on their interactions.
-(c) Action decoder: Uses the final global embedding to construct a grounded action through a GRU-based decoder sequentially - first selecting an action schema, then iteratively choosing objects for each parameter position until a complete grounded action is formed.
-Learning general policies for classical planning domains requires effectively processing variable-sized states and selecting appropriate actions that generalize across problem instances.
-We introduce GABAR (Graph Attention-based Decoder for Action Ranking), a novel architecture that directly learns to rank and select actions rather than learning value functions.
-GABAR consists of four key components: (1) a graph-based state representation that explicitly captures grounded action information, (2) a neural encoder that processes this rich representation, and (3) a GRU-based decoder that sequentially constructs complete actions.
-This section details each component and explains how they work together to enable effective action selection.
-System Overview
-Given a planning instance, I in PDDL format (Fox and Long 2003), along with the set of ground actions, GABAR operates by first converting this to a graph structure that makes explicit the relationships between objects, predicates, and potential actions.
-This graph is then processed through our neural architecture to rank actions(as described in Fig 1).
-Then, the highest-ranked applicable action is executed to reach the next state I’.
-This process is repeated until a goal state is reached.
-To ensure the execution terminates, the system maintains a history of visited states and avoids actions that would lead to previously visited states.
-The execution continues until either reaching a goal state or terminating if no unvisited successor states are available or if the maximum execution length (1000 in our experiments) is exceeded.
-Graph Representation
-We introduce a novel graph representation for classical planning tasks that captures the structural relationships between objects, predicates, and actions and the semantic information needed for learning action ranking effectively.
-Our representation $G=(V,E,X,R)$ consists of a set of nodes $V$, edges $E$, node features $X$, and edge features $R$.
-The node set $V=O\cup P\cup A\cup\{g\}$ where $O$, $P$,and $A$ represent sets of domain objects, grounded predicates, and action schemas respectively and $g$ is a global node that aggregates graph-level information.
-The edge set $E=E_{\text{pred}}\cup E_{\text{act}}$, where $E_{\text{pred}}$ is the set of edges between predicates and their argument objects and similarly $E_{\text{act}}$ is the set of edges between action schemas and their argument objects.
-Node Features
-The node feature function $X:V\rightarrow\mathbb{R}^{d}$ maps each node to a feature vector that encodes type and semantic information.
-The feature vector is constructed by concatenating several one-hot encoded segments:
-For any node $v\in V$: $X(v)=[X_{\text{type}}(v)\parallel X_{\text{act}}(v)\parallel X_{\text{pred}}(v)\parallel X_{\text{obj}}(v)]$, where:
+(b) GNN encoder: Processes the graph through $L$ rounds of message passing where edge, node, and global representations are sequentially updated (c) Action decoder: Uses the final global embedding to construct a grounded action through a GRU-based decoder sequentially - first selecting an action schema, then iteratively choosing objects for each parameter position until a complete grounded action is formed.
+3.2 Graph Representation
+We introduce a novel graph representation (shown in Fig 2) for classical planning tasks that captures the structural relationships between objects, predicates, actions, and the semantic information needed to effectively learn action ranking.
+Our representation $G=(V,E,X,R)$ consists of a set of nodes $V$, edges $E$, node features $X$ and edge features $R$.
+The node set $V=O\cup P\cup A\cup\{g\}$ where $O$, $P$,and AA represent sets of domain objects, grounded predicates (predicates instantiated in current and goal states), and action schemas, respectively, and $g$ is a global node that aggregates graph-level information.
+The edge set $E=E_{\text{pred}}\cup E_{\text{act}}$, where
+$E_{\text{pred}}$ is the set of edges between predicates and their argument objects and similarly $E_{\text{act}}$ is the set of edges between action schemas and their argument objects.
+Node Features: The node feature function $X:V\rightarrow\mathbb{R}^{d}$ maps each node to a feature vector that encodes type and semantic information.
+The feature vector is constructed by concatenating several one-hot encoded segments.
+For any node $v\in V$, $X(v)=[X_{\text{type}}(v)\parallel X_{\text{act}}(v)\parallel X_{\text{pred}}(v)\parallel X_{\text{obj}}(v)]$, where,
+Figure 2: Example graph construction for a simplified blocksworld problem with only onon and clearclear predicates.
+The left side shows the starting and goal states.
+The $O$ nodes are the object nodes.
+In the start state, $O3$ is on $O2$ which is on the table, and $O1$ is on the table.
+In the goal state, $O3$ is on $O1$ which is on the table, and $O2$ is on the table.
+The right side shows the constructed graph with action nodes (blue), object nodes (yellow), and predicate nodes (red).
+The “Pick-up” action connects to object $O1$, while the “Unstack” action connects to objects $O2$ and $O3$.
+Predicate nodes show the current state (“Clear” for $O1$ and $O3$, “On (O3,O2)”) and goal state (“G-ON(O3,O1)”).
+Blue edges represent action-object connections, while red edges represent predicate-object relationships.
   • $X_{\text{type}}\in\{0,1\}^{3}$: One-hot encoding of node type (object, predicate, or action)
   • $X_{\text{act}}\in\{0,1\}^{|A|}$: One-hot encoding of action type (if $v$ is an action node)
-  • $X_{\text{pred}}\in\{0,1\}^{2|P|}$: Encoding for predicates, where first $|P|$ bits indicate predicate type and next $|P|$ bits indicate goal predicates ((if $v$ is a predicate node))
-  • $X_{\text{obj}}\in\{0,1\}^{|T|}$: One-hot encoding of object type (if $v$ is an object node), where $T$ is the set of object types
-Edge Features
-The edge feature function $R:E\rightarrow\mathbb{R}^{k}$ maps each edge to a feature vector encoding edge type and role information:
-For any edge $e\in E$: $R(e)=[R_{\text{type}}(e)\parallel R_{\text{pred}}(e)\parallel R_{\text{act}}(e)]$, where:
+  • $X_{\text{pred}}\in\{0,1\}^{2|P|}$: Encoding for predicates, where first $|P|$ bits indicate predicate type and next $|P|$ bits indicate goal predicates (if $v$ is a predicate node)
+  • $X_{\text{obj}}\in\{0,1\}^{|T|}$: One-hot encoding of the object type (if $v$ is an object node). $T$ is the set of object types
+Edge Features: The edge feature function $R:E\rightarrow\mathbb{R}^{k}$ maps each edge to a feature vector that encodes the type of edge and the information about the role.
+For any edge $e\in E$, $R(e)=[R_{\text{type}}(e)\parallel R_{\text{pred}}(e)\parallel R_{\text{act}}(e)]$, where,
   • $R_{\text{type}}\in\{0,1\}^{2}$: One-hot encoding of edge type (predicate-object or action-object)
@@
     • One-hot encoding of parameter position in action schema ($m$ bits)
-    • Binary vector indicating which predicates are satisfied by the object in the grounded action ($|P|$ bits)
-Global Features
-The global node $g$ is initialized with a zero vector in $\mathbb{R}^{h}$ where $h$ is the chosen hidden representation dimension.
+    • Binary vector indicating which predicates are satisfied by the object in grounded action ($|P|$ bits)
+Global Features: The global node $g$ is initialized with a zero vector in $\mathbb{R}^{h}$ where $h$ is the chosen hidden representation dimension.
 This node can be used to aggregate and propagate graph-level information during message passing.
+This global node is essential in passing information at each round from nodes that are far apart.
+Hence, while nodes get their location information from neighboring edges, they also receive some global context at each round of the GNN ensuring that the number of GNN rounds does not become a bottleneck in passing information in large graphs.
+This is important since, as problems get larger the graph size grows, but the number of GNN rounds stays constant.
+The global node ensures all nodes and edges in the graph have some global context at each step.
 This graph representation captures both the structural and semantic information necessary for learning planning heuristics while maintaining a bounded feature dimension independent of problem size.
 The node features encode type and semantic information, while the edge features capture the relationships between objects, predicates, and actions in the planning domain.
-Neural Architecture
-Our neural architecture processes this graph representation through multiple components designed to handle the challenges of processing variable-sized inputs, capturing long-range dependencies between objects and actions, and making sequential decisions to construct complete actions.
+3.3 Neural Architecture
+Our neural architecture processes this graph representation through multiple components designed to handle the challenges of processing variable-sized inputs, capturing long range dependencies between objects and actions, and making sequential decisions to select actions and their arguments as shown in algorithm 1.
 We detail each component below:
-Graph Neural Network Encoder
-Graph Neural Networks (GNNs) (Scarselli et al.
-2008) are particularly well-suited for encoding states in planning problems as they can naturally process relational structures while being invariant to permutations and handling varying input sizes.
+3.3.1 GNN Encoder
+Graph Neural Networks (GNNs) (gnn_base, ) are particularly well-suited for encoding states in planning problems as they can naturally process relational structures while being invariant to permutations and handling varying input sizes (gnn_powerful, ).
 This allows them to learn patterns that generalize across different problem instances within the same domain, regardless of the number of objects involved.
 The key insight is that planning states are inherently relational - objects interact through predicates and actions - and GNNs can capture these relationships through message passing between nodes and edges.
-Our GNN encoder processes the input graph through $L$ layers of message passing between nodes, edges, and a global node. At each layer $l$, the updates proceed as follows:
-First, edge embeddings are updated based on their incident nodes and the global context:
-$\mathbf{e}^{l+1}_{ij}=\phi_{e}([\mathbf{e}^{l}_{ij};\mathbf{v}^{l}_{i};\mathbf{v}^{l}_{j};\mathbf{g}^{l}])$ (1)
-where $\mathbf{e}^{l}_{ij}$ is the embedding of edge $(i,j)$ at layer $l$, $\mathbf{v}^{l}_{i}$ and $\mathbf{v}^{l}_{j}$ are the embeddings of its incident nodes, $\mathbf{g}^{l}$ is the global node embedding, [;][;][ ; ] denotes concatenation, and $\phi_{e}$ is a learnable neural network.
-Next, node embeddings are updated using information from their connected edges and the global context:
-$\mathbf{v}^{l+1}_{i}=\phi_{v}([\mathbf{v}^{l}_{i};\text{AGG}(\{\mathbf{e}^{l+1}_{ij}|j\in\mathcal{N}(i)\});\mathbf{g}^{l}])$ (2)
-Where AGG is a permutation-invariant aggregation function (we use attention weighted sum), $\mathcal{N}(i)$ denotes the neighbors of node $i$, and $\phi_{h}$ is a learnable network.
-Finally, the global node is updated by aggregating information from all nodes and edges:
-$\mathbf{g}^{l+1}=\phi_{g}([\mathbf{g}^{l};\text{AGG}(\{\mathbf{v}^{l+1}_{i}|i\in\mathcal{V}\});\text{AGG}(\{\mathbf{e}^{l+1}_{ij}|(i,j)\in\mathcal{E}\})])$ (3)
-where $\phi_{g}$ is a learnable network that combines the previous global state with aggregated node and edge information.
+The GNN encoder (Algorithm 1, left side) transforms the input graph into learned embeddings.
+Initialization (line 1): We initialize node features, edge features, and a global feature vector from the input graph.
+Message Passing (lines 3-9): For $L$ iterations, the encoder performs the following three update steps.
+1) Edge Update (line 5): Each edge updates its representation by combining features from its connected elements.
+2) Node Update (lines 7-8): Each node aggregates information from its neighboring edges along with global context.
+3) Global Update (lines 9-10): The global vector is updated by aggregating information from all nodes and edges.
+The aggregation function (AGG) employs attention mechanisms to weight different elements according to their relevance.
 This architecture makes several key design choices motivated by the planning domain.
@@
 Second, we include a global node that can rapidly propagate information across the graph.
-This is particularly important as planning problems scale up - without the global node, information would need to flow through many message-passing steps to reach distant parts of the graph.
+This is important because without the global node, information would need to flow through many message-passing steps to reach distant parts of the graph.
 The global node acts as a shortcut, allowing the model to maintain a comprehensive view of the planning state even as the number of objects and relations grows.
 After $L$ rounds of message passing have been performed, the final global node embedding $g^{l+1}$ captures the relevant planning context needed for action selection.
-GRU-based Action Decoder
-One of the key novelties of our architecture is a GRU-based action decoder that supports actions with an arbitrary number of parameters.
-The decoder transforms the final global graph representation $g_{l+1}$ into a sequence of parameter decisions that fully specify an action.
-This sequential process is managed by a Gated Recurrent Unit (GRU) that maintains a hidden state $h_{i}$ representing the partial action construction:
-$v_{1}$ $=\text{GRU}(g_{v+1},0)$ // Initialize with global state
 […]
-The GRU architecture allows the decoder to maintain relevant context while making a sequence of interdependent decisions.
-This is crucial for ensuring that selected objects form valid and effective action instantiations.
-During training, action selection proceeds greedily in two phases: First, the decoder selects an action schema using the initial hidden state:
-$\text{score}(a)=\text{MLP}([v_{1};a])\quad\text{for each action schema }a$
-$a^{*}=\arg\max_{a\in A}\text{softmax}(\text{score}(a))$
-Then, it iteratively selects the action parameters:
-$\text{score}(o)=\text{MLP}([v_{i};o])\quad\text{for each candidate object }o$
-$o^{*}=\arg\max_{o\in O}\text{softmax}(\text{score}(o))$
-$v_{i+1}=\text{GRU}(v_{i},\text{embed}(o^{*}))$
+1: Initialize: $\{\mathbf{v}^{0}_{i}\}_{i\in\mathcal{V}}$, $\{\mathbf{e}^{0}_{ij}\}_{(i,j)\in\mathcal{E}}$, $\mathbf{g}^{0}$
+[…]
+[…]
+1: $h_{1}=\text{GRU}(\mathbf{g}^{L},\mathbf{0})$
+2: // Select action schema 3: $\text{score}(a)=\text{MLP}(h_{1}\odot\mathbf{v}^{L}_{a})$ for each $a\in A$ 4: $a^{*}=\arg\max_{a\in A}\text{softmax}(\text{score}(a))$ 5: $h_{2}=\text{GRU}(h_{1},\mathbf{v}^{L}_{a^{*}})$ 6: // Select parameters 7: for $i=1$ to required parameter count do 8:   $\text{score}(o)=\text{MLP}(h_{i+1}\odot\mathbf{v}^{L}_{o})$ for each $o\in O$ 9:   $o^{*}_{i}=\arg\max_{o\in O}\text{softmax}(\text{score}(o))$ 10:   $h_{i+2}=\text{GRU}(h_{i+1},\mathbf{v}^{L}_{o^{*}_{i}})$ 11: return $(a^{*},o^{*}_{1},o^{*}_{2},\ldots,o^{*}_{n})$
+Algorithm 1 Graph Attention-Based Action Ranking (GABAR)
+3.3.2 GRU Decoder
+The GRU-based decoder (Algorithm 1, right side) uses the encoded graph to select an action and its parameters.
+Initialization (line 1): The decoder initializes its hidden state using the global graph embedding and a zero vector.
+Action Schema Selection (lines 3-5): The decoder computes a score for each potential action schema (line 3).
+It then Selects the highest-scoring action (line 4) and updates the hidden state with the selected action’s features (line 5).
+Parameter Selection (lines 7-10): For each required parameter position, scores are computed for all objects using the current hidden state (line 8).
+The highest-scoring object is selected as the parameter (line 9).
+The hidden state is updated with the selected object’s features (line 10).
+Output (line 11): The decoder returns the complete action grounding.
+This autoregressive approach ensures that each parameter selection is conditioned on the graph structure, the selected action schema, and all previously selected parameters.
+By updating the GRU hidden state with each selection, the model captures dependencies between different components of the action.
+We repeat this greedily until all action parameters have been extracted.
 Since greedy parameter selection is often too myopic, we employ beam search to explore multiple choices in parallel.
@@
 The final output is a ranked list of $k$ action groundings $(a,o_{1},\ldots,o_{n})$ along with their accumulated scores.
-This aligns with our goal of learning to rank actions rather than just selecting individual best choices - while training optimizes for selecting the optimal action, the learned model provides a ranking of top $k$ grounded actions during execution.
-The planner can then use this ranking to make more informed decisions, such as incorporating additional criteria such as cycle avoidance.
-Data Generation and Training
+This process of GRU-based decoding helps generalize the framework to handle actions with any arity, as we can extract as many objects as necessary for the action selected in the first step of the decoding.
+Hence, at any given state, we can use the decoder to extract the ranking over the set of actions with different arity all at once.
+This aligns with our goal of learning to rank fully grounded actions rather than independently optimizing each argument (extended beam search version of the decoder that maintains multiple candidate action groundings in parallel is provided as algorithm 2 in the Appendix).
+While training optimizes for selecting the optimal action, the learned model provides a ranking of the top $k$ grounded actions during execution.
+The planner then selects the first legal action in the current state according to the ranking.
+It also uses the ranking to avoid cycles - if an action leads to a state previously visited, then the next best action is considered.
+3.4 Data Generation and Training
 For each planning domain, we generate training data by solving a set of small problem instances using an optimal planner.
@@
   • $y_{o}=\{y_{o1},...,y_{ok}\}$: A sequence of $k$ one-hot vectors over the object space, where $k$ is the maximum number of parameters any action can take, indicating the correct objects for each parameter position
-For action schema selection: the model needs to learn to assign the highest score to the correct action schema among all possible schemas.
-For object selection: for each parameter position, the model needs to learn to assign the highest score to the correct object among all candidate objects.
+For action schema selection, the model needs to learn to assign the highest score to the correct action schema among all possible schemas.
+For each parameter position, the model needs to learn to assign the highest score to the correct object among all candidate objects.
 This is done using the following loss function.
-Loss Function Given a training instance $(G,y_{a},y_{o})$, GABAR computes action scores $s_{a}$ for all possible action schemas and object scores $s_{o_{i}}$ for each parameter position $i$.
+Loss Function.
+Given a training instance $(G,y_{a},y_{o})$, GABAR computes action scores $s_{a}$ for all possible action schemas and object scores $s_{o_{i}}$ for each parameter position $i$.
 The total loss is computed as $L=L_{action}+L_{objects}$, where, $L_{objects}$ is the sum of cross-entropy losses between $\text{softmax}(s_{o_{i}})$ and $y_{o_{i}}$ for each parameter position $i$.
-Training Procedure We train the model using the Adam optimizer with a learning rate of $0.0005$, $9$ rounds of GNN, and batch size of $16$.
+Training Procedure.
+For all domains, we train the model using the Adam optimizer with a learning rate of $0.0005$, 99 rounds of GNN, and batch size of $16$, hidden dimensionality of $64$.
 Training proceeds for a maximum of 500 epochs, and we select the model checkpoint that achieves the lowest loss on the validation set for evaluation.
-Experiments
-Table 1: Performance of both GABAR and GABAR-G methods across domains(number of problems shown in parenthesis(#)).
-Coverage indicates the percentage of test instances solved within 1000 steps.
-Plan lengths are reported only for solved instances.
-GABAR GABAR-G
-Domain Cov.(%) $\uparrow$ Mean PL $\downarrow$ Median PL $\downarrow$ PQR $\uparrow$ Cov.(%) $\uparrow$ Mean PL $\downarrow$ Median PL $\downarrow$ PQR $\uparrow$
-Blocks World (100) 100 53.1 46 1.42 84 52.6 54.7 1.1
-Gripper (100) 100 147.1 133 0.99 89 146.7 137 0.95
-Miconic (50) 100 181.7 181 0.97 75 192.4 189 0.87
-Logistics (50) 76 155.2 132 0.65 12 238.5 220 0.33
-Visitall (50) 88 207.9 190 1.07 63 205.7 183 0.86
-Grid (50) 96 72.1 62 0.81 68 70.2 68 0.73
+It takes between 1-2 hours to train a model for each domain on an RTX 3080.
+4 Experiments
 We evaluate GABAR’s performance across a diverse set of classical planning domains.
 Our experiments aim to assess both the quality of learned policies and their ability to generalize to significantly larger problems than those in training.
-Domains. We selected six standard planning domains that present different types of structural complexity and scaling dimensions.
+We selected eight standard planning domains that present different types of structural complexity and scaling dimensions.
+The domains are : BlocksWorld, Gripper, Miconic, Spanner, Logistics, Rovers, Grid and Visitall (More details in appendix A.1).
+We divide the test set for each domain into 3 separate subsets, easy, medium, and hard with increasing difficulty with problem sizes as defined in table 1 along with the training and validation dataset sizes.
+Each test subset has 100 problems.
+In contrast, the training dataset consists of problems smaller and simpler than the ones found in the easy subset.
+Table 1: Distribution of problem sizes across train, validation, and test datasets.
+The ranges indicate the number of objects/variables defining each domain’s problem complexity.* In Visitall and Grid, size refers to number of cells
+domain Train Val Test domain Train Val Test
+easy med hard easy med hard
+Blocks $[6,9]$ $[10]$ $[11,20]$ $[21,30]$ $[31,40]$ Rovers $[3,9]$ $[10]$ $[11,30]$ $[31,50]$ $[51,70]$
+Gripper $[5,15]$ $[17]$ $[20,40]$ $[41,60]$ $[61,100]$ Visitall* $[9,36]$ $[49]$ $[50,100]$ $[101,200]$ $[201,400]$
+Miconic $[1,9]$ $[10]$ $[20,40]$ $[41,60]$ $[61,100]$ Grid* $[25,49]$ $[63]$ $[64,100]$ $[100,125]$ $[125,154]$
+Spanner $[2,9]$ $[10]$ $[11,20]$ $[21,30]$ $[31,40]$ Logistics $[4,7]$ $[8]$ $[15,20]$ $[21,25]$ $[26,30]$
+We evaluate GABAR against three state-of-the-art approaches for learning generalized policies: GPL, ASNets, and GRAPL.
+Each represents a different architectural paradigm for tackling generalized planning.
+4.1 Baseline Approaches
+GPL (Generalized Policy Learning) unsupervised_gnn learns value functions over states using GNNs.
+It selects actions by identifying unvisited states with the lowest estimated cost-to-goal.
+GPL attempts to learn a globally consistent value function by minimizing the regularized Bellman error $|V(s)-(1+\min_{s^{\prime}\in N(s)}V(s^{\prime}))|+\max\{0,V^{*}(s)-V(s)\}+\max\{0,V(s)-\delta V^{*}(s)\}$ where $s$ is state and $V(s)$ is value function defined as cost to goal.
+ASNets (Action Schema Networks) toyer2020asnets employs a neural network with alternating action and proposition layers.
+Each layer contains modules that connect only to related modules in adjacent layers.
+ASNets shares weights across modules of the same action schema or predicate, enabling generalization across problems of varying sizes but limiting its reasoning to a fixed-depth receptive field.
+GRAPL (Graph Relational Action Ranking for Policy Learning) karia2021learning learns to rank actions using canonical abstractions where objects with identical properties are grouped together.
+Its neural network predicts both action probabilities and plan length estimates but lacks the explicit action-object graph structure and conditional decoding mechanisms of our approach.
+Ablations: To assess each component’s contribution to GABAR, we conducted ablation experiments with three variants:
+  • GABAR-ACT: Removes action nodes and action-object edges from the graph, using only object and predicate nodes for decision-making.
+  • GABAR-CD: Eliminates conditional decoding by selecting action parameters independently rather than sequentially. This tests the importance of dependencies between action parameters.
+  • GABAR-RANK: Replaces action ranking with a value function learning objective while maintaining our graph representation.
+These ablations isolate our key innovations—action-centric graphs, GRU-based conditional decoding, and action ranking—to measure their individual impact on performance.
+Evaluation Metrics. We evaluate GABAR using two primary metrics:
+  • Coverage (C): The percentage of test instances successfully solved within a 1000-step limit.
+  • Plan Quality Ratio(P): Ratio of plan length produced by Fast Downward (FD) (fast_downward, ) planner to the plan length produced by the learned policy.
+FD is run with fd-lama-first setting.
+We chose this satisficing configuration over optimal planners since optimal planners fail to solve most test problems within a reasonable time.
+While this means we cannot guarantee the optimality of the reference plans, it provides a practical baseline for assessing solution quality.
+High coverage on larger instances demonstrates the model’s ability to learn general action selection strategies, while plan quality ratio reveals whether these strategies produce efficient plans as problems scale up.
+5 Results and Discussions
+Table 2: Coverage (C) and plan quality ratio (P) across different domains and difficulty levels.
+Turqoise color intensity for C values indicates coverage score (Only values above 50% are highlighted: Light: 50-74%, Medium: 75-89%, Dark: 90-100%).
+P values are highlighted in violet with similar thresholds (No highlight: 0-0.49, Light: 0.5-0.74, Medium: 0.75-0.99, High: 1.0+).
+Columns are grouped into Baselines and Ablations.
+Combined rows show averages across all domains.
+(*) only non-zero values from their respective domains were considered
+Baselines Ablations
+GPL ASNets GRAPL GABAR GABAR-ACT GABAR-CD GABAR-RANK
+[…]
+Blocks E 100 1.1 100 1.6 64 0.65 100 1.5 44 0.65 100 0.92 29 0.79
+M 45 0.68 100 1.5 48 0.44 100 1.6 14 0.49 92 0.81 21 0.71
+H 10 0.33 92 1.4 38 0.28 100 1.7 4 0.35 81 0.80 9 0.61
+Miconic E 97 0.97 100 1.0 68 0.56 100 1.0 35 0.55 97 0.88 42 0.67
+M 37 0.56 100 0.98 65 0.54 100 0.97 18 0.33 94 0.86 29 0.37
+H 19 0.29 90 0.92 60 0.49 100 0.95 2 0.27 88 0.83 16 0.29
+Spanner E 73 1.1 78 0.86 22 0.65 94 1.1 31 0.65 87 0.98 57 0.82
+M 42 0.56 60 0.69 5 0.55 93 0.99 11 0.27 81 0.93 42 0.77
+H 3 0.18 42 0.61 0 - 89 0.91 0 - 62 0.79 12 0.45
+Gripper E 100 1.0 78 0.98 26 0.95 100 1.1 31 0.56 95 1.0 55 0.58
+M 56 0.85 54 0.91 12 0.67 100 0.99 23 0.40 92 0.93 43 0.41
+H 21 0.74 42 0.88 0 - 100 0.96 9 0.28 87 0.86 21 0.33
+Visitall E 69 1.3 94 0.96 92 1.1 93 1.1 72 1.2 91 1.1 52 0.64
+M 15 0.76 86 0.93 88 1.0 91 1.0 64 0.93 89 1.1 46 0.56
+H 0 0 64 0.81 78 0.99 88 1.1 44 0.67 83 1.2 39 0.54
+Grid E 74 0.89 52 0.81 20 0.38 100 0.91 21 0.56 79 0.87 17 0.54
+M 17 0.61 45 0.66 3 0.28 97 0.85 8 0.46 71 0.65 12 0.28
+H 0 0 21 0.60 0 - 92 0.74 0 - 54 0.53 0 -
+Logistics E 56 0.61 39 0.71 32 0.81 90 0.75 12 0.64 31 0.86 41 0.65
+M 7 0.21 22 0.55 9 0.45 76 0.65 3 0.49 25 0.54 21 0.49
+H 0 0 4 0.39 0 - 71 0.59 0 - 6 0.35 0 -
+Rovers E 64 0.99 67 0.96 21 0.35 87 1.0 22 0.75 44 0.81 33 0.67
+M 9 0.32 56 0.87 5 0.19 82 0.96 6 0.66 37 0.63 9 0.56
+H 0 0 31 0.64 0 - 77 0.97 0 - 19 0.57 0 -
+Combined E 79.1 0.98 76 0.98 43.5 0.67 95.5 1.0 33.5 0.69 78 0.93 40.2 0.67
+M 28.5 0.56 65.4 0.88 29.3 0.51* 92.2 1.0 18.4 0.50 72.7 0.80 27.8 0.51
+H 6.5 0.39* 48.5 0.78 22.1 0.58* 89.2 0.99 7.4 0.39* 60 0.73 12.1 0.44*
+Table 2 shows the performance of GABAR, its ablations, and the baseline approaches across all domains.
+The results demonstrate significant differences in generalization capabilities and solution quality.
+5.1 Generalization Performance
+GABAR generalizes remarkably well across all domains, with coverage dropping minimally from 95.5% on easy problems to 89.5% on hard problems.
+This generalization is particularly evident in domains like Visitall, where GABAR solves problems 8 times larger than those in training.
+On BlocksWorld, Gripper, and Miconic, GABAR achieves 100% success rate at all difficulty levels.
+Even in complex domains requiring multi-relation features (Grid, Logistics) or higher arity predicates (Rovers), GABAR maintains strong performance with success rates of 92%, 62%, and 77% respectively on hard instances.
+This demonstrates that our graph structure effectively encodes critical problem information, and the decoder successfully captures relationships between actions and their parameters.
+We now analyze GABAR against baselines and relevant ablation results, highlighting how our design choices contribute to performance improvements.
+Action Ranking vs.
+Value Function Learning: GPL uses GNNs to learn value functions over states, aiming to produce globally consistent estimates of cost-to-goal unsupervised_gnn .
+As problems grow larger, maintaining these consistent estimates becomes significantly more challenging.
+GPL’s coverage drops to 0% on hard instances of Logistics and Visitall where GABAR maintains 71% and 88% coverage.
+Our ablation GABAR-RANK, which replaces action ranking with value function learning while keeping our graph structure intact, shows a similar pattern—coverage decreases from 89.2% to just 12.1% on hard problems.
+This parallel between GPL and GABAR-RANK confirms that directly ranking actions is more effective than learning value functions.
+Conditional vs.
+Non-conditional Decoding: GRAPL also learns action rankings but uses an abstract state representation without explicit modeling of action parameter dependencies karia2021learning .
+While conceptually similar to our approach, GRAPL achieves only moderate performance on easy instances (43.5% coverage) and struggles on harder problems (22.1% coverage) compared to GABAR’s 95.5% and 89.2%.
+Our ablation GABAR-CD, which removes our conditional decoding mechanism, shows a similar degradation—coverage drops from 89.2% to 60.0% on hard problems overall, with particularly dramatic decreases in complex domains like Logistics (from 62% to 21%).
+This comparison highlights the critical importance of our conditional GRU-based decoder for handling complex dependencies in domains where objects participate in multiple predicates simultaneously (e.g., a package being in both a city and a vehicle).
+The decoder learns relationships between parameter selections by conditioning each object choice on previously selected objects, enabling the model to capture complex inter-parameter relationships that GRAPL cannot.
+Action-Centric Graph Structure: ASNets employs alternating action and proposition layers with a fixed-depth receptive field toyer2020asnets , limiting its ability to reason about long chains of dependencies.
+While ASNets performs reasonably on medium-difficulty problems, it generalizes poorly to larger instances—achieving only 21%, 4%, and 31% coverage on hard instances of Grid, Logistics, and Rovers, compared to GABAR’s 92%, 62%, and 77%.
+Although we don’t have a perfect ablation parallel for ASNets, our GABAR-ACT ablation (removing action nodes) shows catastrophic performance drops from 95.5% to 33.5% on easy problems and from 89.2% to just 7.4% on hard problems.
+This suggests that explicit action-object relationship modeling in the graph structure is critical for robust generalization.
+These comparisons demonstrate that GABAR’s superior performance stems from the synergistic combination of three key ideas: (1) an action-centric graph representation that captures relationships between actions and objects, (2) conditional decoding that models dependencies between action parameters, and (3) direct action ranking that avoids the challenges of learning globally consistent value functions.
+The ablation studies confirm that each component addresses these limitations present in current approaches to generalized planning.
+6 Conclusion and Future Work
+We presented GABAR, a novel graph-based architecture for learning generalized policies in classical planning through action ranking.
+Our key contributions include (1) an action-centric graph representation that explicitly captures action-object relationships, (2) a GNN architecture augmented with global nodes and GRUs for effective information propagation, and (3) a sequential decoder that learns to construct complete grounded actions.
+First, GABAR shows strong generalization capabilities, successfully solving problems significantly larger than those used in training.
+The system maintains high coverage (89%) even on the hardest test instances that are up to 8 times larger than training problems.
+This generalization is particularly significant in complex domains like logistics and grid that require coordinating multiple entities.
+Second, our comparison with GPL demonstrates that learning to rank actions is more generalizable and sample-efficient than learning value functions.
+GABAR achieves superior performance using only 25% of the training data required by GPL while maintaining better coverage and plan quality across all difficulty levels.
+Third, our ablation study validates the importance of the global node in facilitating long-range information sharing.
+Without this architectural component, performance degrades significantly on larger problems, particularly in domains requiring complex coordination.
+The primary challenge for the proposed method is the graph representation’s growth rate, particularly for domains with high-arity actions or predicates.
+Future work could explore more compact representations while maintaining expressiveness and investigate ways of pruning irrelevant actions in the graph.
+Other future work could also include expanding the representation to handle uncertainty and solve more real-world problems such as object rearrangement.
+7 Acknowledgments
+The authors acknowledge the support of Army Research Office under grant W911NF2210251
+Appendix A Appendix
+A.1 Domain descriptions
 Blocks World involves manipulating blocks to achieve specific tower configurations.
@@
 Logistics involves transporting packages between locations using trucks (for intra-city transport) and airplanes (for inter-city transport).
-The domain scales with both the number of cities (4-8 for training/validation, 15-30 for testing) and packages (3-10 for training/validation, 9-24 for testing).
+The domain scales with both the number of cities (4-8 for training/validation, 15-30 for testing) and packages (3-9 for training/validation, 10-24 for testing).
 Visitall requires an agent to visit all cells in a grid.
 The domain scales with grid size (9-49 cells for training/validation, up to 400 cells for testing - testing problems 8 times larger than the training dataset).
-Grid involves navigating through a grid where certain doors are locked and require specific keys to open.
-The number of locks (3) and keys (5) remain the same across training and testing while varying the size of the grid (7$\times$9 for training/validation to 11$\times$14 for testing - a 150% increase in cells).
-Evaluation Metrics. We evaluate GABAR using three primary metrics:
-  • Coverage (Cov): The percentage of test instances successfully solved within a 1000-step limit.
-  • Plan Length (PL): Measured through both mean and median plan lengths across solved instances.
-  • Plan Quality Ratio(PQR): Ratio of plan length produced by Fast Downward (FD) planner to the plan length produced by the learned policy.
-FD is run with fd-lama-first setting.
-We chose this satisficing configuration over optimal planners since optimal planners fail to solve most test problems within a reasonable time.
-While this means we cannot guarantee the optimality of the reference plans, it provides a practical baseline for assessing solution quality across our test suite.
-High coverage on larger instances demonstrates the model’s ability to learn robust action selection strategies, while plan length and quality metrics reveal whether these strategies remain efficient as problems scale up.
-Ablation: To understand the importance of different architectural components, we conduct an ablation study focusing on the design choice of using global nodes in the graph (GABAR-G).
-This variant of GABAR is without the global node, which helps assess its role in information propagation and helps quantify how the global node affects coverage, plan quality, and length.
-Results and Discussions
-Table 1 shows the results of our learning system across $6$ domains.
-We can see that the proposed method (OURS) generalizes very well on three domains - Blocksworld, Gripper, and Miconic.
-We can also see that our method generalizes well in the other three domains.
-This result, while not as good as others, shows an important capability of the proposed system - the ability to handle complex domains.
-Both Grid and Logistics are more complex domains (Ståhlberg, Bonet, and Geffner 2022b), as expressing properties of the environment requires multiple relations (Ex, the feature expressing that a package is in a city while possibly within a vehicle needs multiple relations to express).
-We can see that the proposed method can learn these complex features that are required for solving the planning problem as well as generalize to larger problems in these settings (success rate of 76% and 96% in them).
-While the success rate of Visitall is not as impressive as Blocksworld, the generalization it achieves is impressive - it is solving problems up to eight times larger than what it trained on with an 88% success rate.
-The ablation study comparing GABAR with and without the global node (GABAR vs. GABAR-G) reveals that removing the global node leads to degradation in performance.
-In domains that primarily require local reasoning, like Blocks World, the impact is moderate but still notable - coverage drops from 100% to 84%, and the plan quality ratio decreases from 1.42 to 1.1.
-This suggests that while local messages passing between objects and actions can capture basic patterns, the global node helps maintain a broader context.
-This degradation is even more dramatic in Logistics, where coverage plummets from 76% to just 12%.
-The stark difference can be attributed to the need for long-range planning, as logistics requires orchestrating multiple vehicles across different cities.
-Plan quality also suffers significantly without the global node.
-In Logistics, the plan quality ratio drops from 0.65 to 0.33, indicating that solutions become nearly twice as inefficient.
+[…]
+Spanner requires tightening nuts at one end of a corridor with spanners collected along the way.
+Number of locations varies from 10-20 in training and testing, the number of spanners varies from 2-10 in training and 11 to 40 in testing.
+Rovers requires multiple rovers equipped with different capabilities to perform experiments and send results back to the lander.
+Training and validation instances use 2-5 rovers and 3-10 waypoints; those for testing have 3-7 rovers and 11-70 waypoints.
+Table 3 contains information about number of datapoints used for each domain  [^1] Code : https://anonymous.4open.science/r/ltp-4741/ .
+Dataset submitted as supplementary material.
+All problems were generated using openly available PDDL-generators [17].
+Table 3: Number of training datapoints per domain
+Domain Training Datapoints
+Blocksworld 3,348
+Gripper 6,287
+Miconic 4,458
+Spanner 3,801
+Logistics 6,556
+Rovers 3,874
+VisitAll 3,654
+Grid 6,880
+A.2 Ablations
+We present more ablations on our method in this section to further analyze the contribution of different components in our GABAR architecture. The two configurations tested are:
+  • GABAR-G: An ablated version of GABAR that removes the global node from the graph neural network.
+Without the global node, the model lacks a centralized representation that aggregates information across the entire state, potentially limiting its ability to reason about complex relationships spanning multiple objects.
+  • GABAR-ACT_CD: A severely limited version that removes both the action representation from the graph and the conditional decoding mechanism.
+This baseline lacks the capability to explicitly represent actions in the state graph and cannot condition parameter selection on previously chosen actions, effectively testing the importance of both structured action representation and sequential decision-making.
+Results Discussion
+The results from Table 4 clearly demonstrate the effectiveness of our complete GABAR architecture compared to its ablated counterparts:
+Overall Performance: GABAR significantly outperforms both ablated versions across all domains and difficulty levels.
+With average coverage of 95.5%, 92.2%, and 89.2% for easy, medium, and hard problems respectively, GABAR demonstrates robust generalization capabilities.
+GABAR v/s GABAR-G : The dramatic performance drop from GABAR to GABAR-G (especially on hard problems where coverage falls from 89.2% to 42.5%) highlights the critical importance of the global node in our architecture.
+This component enables effective information aggregation across the entire state graph, substantially improving the model’s ability to handle complex planning scenarios.
 This degradation in quality suggests that the global node plays a crucial role in helping the model learn strategic action selection rather than just locally reasonable choices.
-These findings validate our architectural choice of including a global node in GABAR.
-The global node proves essential for facilitating long-range information sharing that helps coordinate actions across distant entities.
-Conclusion and Future Work
-We presented GABAR, a novel graph-based architecture for learning generalized policies in classical planning through action ranking.
-Our key contributions include (1) an action-centric graph representation that explicitly captures action-object relationships, (2) a GNN architecture augmented with global nodes and GRUs for effective information propagation, and (3) a sequential decoder that learns to construct complete grounded actions.
-Our experimental results demonstrate strong generalization capabilities across multiple planning domains.
-GABAR achieves great coverage on Blocksworld, Gripper, Visitall, and Miconic domains when scaling to problems that are many times larger than those used in training.
-The architecture also shows promising results in more complex domains like Grid and Logistics.
-The ablation study removing global nodes highlights their critical role in enabling effective information propagation.
-The primary scaling challenge lies in the graph representation’s growth rate, particularly for domains with high-arity actions or predicates.
-Future work could explore more compact representations while maintaining expressiveness and investigate ways for pruning irrelevant actions in the graph.
-Acknowledgments
-The authors acknowledge the support of Army Research Office under grant W911NF2210251
+GABAR v/s GABAR-ACT_CD : The extremely poor performance of GABAR-ACT_CD (with coverage below 10% in most domains and near-zero on harder problems) confirms that both action representation within the graph and conditional decoding are essential components.
+Without them, the model fails to capture the relationship between actions and their parameters or to maintain consistency in sequential decisions.
+Plan Quality: Beyond just solving more problems, GABAR consistently generates higher-quality plans (as measured by PQR) compared to its ablated counterparts.
+This indicates that the full model not only finds solutions but discovers more efficient paths to the goal.
+Table 4: Performance across domains and difficulty levels.
+Coverage (C) indicates percentage of test instances solved within 1000 steps.
+Plan Quality Ratio (P) reported only for solved instances.
+Turqoise color intensity for C values indicates coverage score (Only values above 50% are highlighted: Light: 50-74%, Medium: 75-89%, Dark: 90-100%).
+P values are highlighted in violet with similar thresholds (No highlight: 0-0.49, Light: 0.5-0.74, Medium: 0.75-0.99, High: 1.0+).
+Combined rows show averages across all domains.
+(*) only non-zero values from respective domains considered
+ACT_CD GABAR GABAR-G
+[…]
+Blocks E 11 0.33 100 1.5 100 1.2
+M 8 0.31 100 1.6 88 1.2
+H 0 - 100 1.7 32 0.70
+Miconic E 8 0.32 100 1.0 95 0.95
+M 3 0.27 100 0.97 70 0.87
+H 0 - 100 0.95 56 0.73
+Spanner E 12 0.43 94 1.1 82 0.93
+M 1 0.24 93 0.99 74 0.81
+H 0 - 89 0.91 54 0.62
+Gripper E 15 0.58 100 1.1 95 1.0
+M 5 0.34 100 0.99 92 0.95
+H 0 - 100 0.96 77 0.90
+Visitall E 23 0.58 93 1.1 77 0.96
+M 19 0.55 91 1.0 66 0.83
+H 16 0.48 88 1.1 29 0.51
+Grid E 0 - 100 0.91 79 0.83
+M 0 - 97 0.85 68 0.62
+H 0 - 92 0.74 33 0.41
+Logistics E 0 - 90 0.75 45 0.43
+M 0 - 76 0.65 22 0.33
+H 0 - 71 0.59 4 0.21
+Rovers E 3 0.45 87 1.0 77 0.85
+M 0 - 82 0.96 72 0.79
+H 0 - 77 0.97 55 0.65
+Combined E 9 0.44* 95.5 1.0 80.2 0.89
+M 4.5 0.34* 92.2 1.0 69.2 0.79
+H 2 0.48* 89.2 0.99 42.5 0.59
+A.3 Beam Search Action Decoder
+The beam search action decoder (Algorithm 2) enhances the basic greedy decoder by maintaining multiple candidate sequences at each decoding step.
+Initialization (line 1): Similar to the greedy approach, the decoder initializes its hidden state using the global graph embedding and a zero vector.
+Action Schema Selection (lines 2-5): Instead of selecting only the highest-scoring action schema, beam search maintains the top-$k$ action schemas.
+For each action schema $a\in A$, a score $s_{a}$ is computed (line 3).
+The $k$ highest-scoring actions form the initial beam $B_{0}$ (lines 4-5), each representing a partial sequence with an action schema, an empty parameter list, the initial hidden state, and the action score.
+Parameter Selection with Beam Search (lines 6-15): During each decoding step, the algorithm expands all sequences in the current beam by considering the top-$k$ object candidates for the next parameter position:
+  • For each sequence in the current beam $B_{t-1}$ (line 9), scores are computed for all objects using the current hidden state (line 10).
+  • The top-$k$ objects $O^{*}$ are identified (line 11).
+  • Each sequence is expanded with each of these top-$k$ objects (lines 12-14), creating $k\times|B_{t-1}|$ new candidate sequences in $B^{\prime}_{t}$.
+  • The hidden state is updated for each new sequence (line 13).
+  • The beam is pruned to keep only the $k$ highest-scoring expanded sequences (line 15), forming the new beam $B_{t}$.
+Output (line 16): After processing all parameter positions, the algorithm returns the set of completed sequences, each forming a fully grounded action.
+In our implementation, we set $k=2$ for the beam search.
+This parameter choice provides a reasonable trade-off between exploration and computational efficiency.
+A beam width of 2 allows the decoder to maintain alternative sequences when faced with uncertainty, which is particularly important for handling invalid actions that might need to be discarded.
+Setting a larger beam width provides better results theoretically but increases computational cost without proportional benefits in our specific domain.
+[…]
+[0.8ex]
+1: $h_{0}=\text{GRU}(\mathbf{g}^{L},\mathbf{0})$
+[0.8ex] 2: // Step 1: Action schema selection
+[0.5ex] 3: $s_{a}=\text{MLP}(h_{0}\odot\mathbf{v}^{L}_{a}),\forall a\in A$
+[0.5ex] 4: $A^{*}=\{a_{1},a_{2},\ldots,a_{k}\}=\underset{a\in A}{\text{top-}k}(s_{a})$
+[0.5ex] 5: $B_{0}=\{(a_{j},[],h_{0},s_{a_{j}})|a_{j}\in A^{*}\}$
+[0.8ex] 6: // Step 2: Parameter selection with beam search
+[0.5ex] 7: for $t=1$ to max_params do
+[0.5ex] 8:   $B^{\prime}_{t}=\emptyset$
+[0.5ex] 9:   for $(a,[o_{1},\ldots,o_{t-1}],h_{t-1},s)\in B_{t-1}$ do
+[0.5ex] 10:    $s_{o}=\text{MLP}(h_{t-1}\odot\mathbf{v}^{L}_{o}),\forall o\in O$
+[0.5ex] 11:    $O^{*}=\{o_{1},o_{2},\ldots,o_{k}\}=\underset{o\in O}{\text{top-}k}(s_{o})$
+[0.5ex] 12:    for $o_{j}\in O^{*}$ do
+[0.5ex] 13:    $h_{t}=\text{GRU}(h_{t-1},\mathbf{v}^{L}_{o_{j}})$
+[0.5ex] 14:    $B^{\prime}_{t}=B^{\prime}_{t}\cup\{(a,[o_{1},\ldots,o_{t-1},o_{j}],h_{t},s+s_{o_{j}})\}$
+[0.5ex] 15:   $B_{t}=\underset{b\in B^{\prime}_{t}}{\text{top-}k}(\text{score}(b))$
+[0.8ex] 16: return $\{(a,[o_{1},o_{2},\ldots,o_{n}])|(a,[o_{1},o_{2},\ldots,o_{n}],h,s)\in B_{\text{max\_params}}\}$
+Algorithm 2 Beam Search Action Decoder